OpenStudy (anonymous):

What is the result of adding the system of equations? 2x + y = 4 3x - y = 6

OpenStudy (anonymous):

You can substitute the y in. 2x+y=4 3x-y=6 Is that how it is set up on your example

OpenStudy (anonymous):

If so, subtract 2x on both sides and get "y" by itself y=4-2x Then substitute where "y" is 3x-(4-2x)=6 Distribute the negative sign, and combine like terms 3x-4+2x=6 5x-4=6 Add 4 to the right side to get "x" by itself, then divide 5 5x=10 5x=10/5 x=2 Answer!